assuming that the sample come from the normal distributions find the margim of error. E given the following





B.n=16,X =50, s=4.2, 95% confidence







1
Expert's answer
2022-06-08T17:09:08-0400

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=15" and the critical value for a two-tailed test is "t_c =2.131449."

"E=t_c\times\dfrac{s}{\sqrt{n}}=2.131449\times \dfrac{4.2}{\sqrt{16}}"

"=2.238"